OpenStudy (anonymous):

<3=<4 m<3=___ m<4=___

OpenStudy (jhannybean):

What is the shape you are given?

OpenStudy (camper4834):

hi it is asking for the measurement of the angle 3 and the angle 4

OpenStudy (camper4834):

do you know how many degrees those are?

OpenStudy (jhannybean):

Better question: How many degrees are there in a circle?

OpenStudy (camper4834):

....why is that a better question

OpenStudy (jhannybean):

If you tell me what you think, I can explain why.

OpenStudy (jhannybean):

Perhaps it will help understand more things about angles and how to find them.

OpenStudy (anonymous):

A contractor wishes to build 9 houses, each different in design. In how many ways can he place these houses on a street if 6 lots are on one side of the street and 3 lots are on the opposite side?

Directrix (directrix):

@Johnyoungman224 Angles 3 and 4 form a striaght angle and therfore sum to 180 degrees. Because the two angles are given to have the same measure, the M<3 + M<4 =180 Let M<3 = x so M<4 = x and x + x = 180 2x = 180 so what is the value of x? That is also the measure of angles 3 and 4.

OpenStudy (jhannybean):

 a circle is 360 degrees. if you divide it up into 4 quadrants, you get 360/4 = 90 . In your drawing you showed that a straight line is cut by a perpendicular bisector, meaning it spits the line AND the angle the line creates directly in half. Therefore you have congruent angles next to eachother that form 90 degrees.